# Scale Where It Matters: Training-Free Localized Scaling for Diffusion Models

### v2

@@ B.3 Localized Resampling Procedure
1:Input image $x_{\mathrm{in}}$ (at t=0t{=}0), prompt $p$, total steps $T$, re-noise step $t_{0}$, integration step $t_{g}$, mask $M$, diffusion model, verifier $V$
@@
5:Global re-noise to $t_{0}$
6:Sample $z_{\text{bg}},z_{\text{mask}}\sim\mathcal{N}(0,I)$
[…]
8:Masked refinement ($t_{0}\rightarrow t_{g}$)
9:for $t=t_{0},\,t_{0}-\Delta t,\,\dots,\,t_{g}$ do
10:   Sample $z_{t}\sim\mathcal{N}(0,I)$
11:   $x_{t-\Delta t}\leftarrow(1-M)\odot(\alpha(t)x_{\text{anchor}}+\sigma(t)z_{t})+M\odot(x_{t}-\Delta t\,\epsilon_{\theta}(x_{t},t;p)+\sigma(t)z_{t})$
12:end for
13:Global integration ($t_{g}\rightarrow 0$)
14:for $t=t_{g},\,t_{g}-\Delta t,\,\dots,\,0$ do
15:   Sample $z_{t}\sim\mathcal{N}(0,I)$
16:   $x_{t-\Delta t}\leftarrow x_{t}-\Delta t\,\epsilon_{\theta}(x_{t},t;p)+\sigma(t)z_{t}$
17:end for
@@
20:return $x_{\mathrm{out}},\,v_{\mathrm{out}}$
21:$T$ is included for scheduler compatibility; discretization uses $\Delta t$.
@@ B.4 Overall Algorithm of LoTTS
1:Prompt $p$, DFS depth D=2D{=}2, global seeds $S$, localized refinements $K$, total steps $T$, re-noise $t_{0}$, integration $t_{g}$, verifier $V$
@@
22:     for $k=1$ to $K$ do
23:        $(x_{\mathrm{out}},v_{\mathrm{out}})\leftarrow\text{LocalizedResample}(x,p,T,t_{0},t_{g},M,V)$
@@ C.1 Mask Generation Examples
Fig. 1 and Fig. 2 provide a detailed visualization of the mask generation process in SD2.1 and FLUX. Across both models, LoTTS exhibits several consistent behaviors that highlight its ability to identify and isolate low-quality regions.
First, the original cross-attention map ($\mathrm{CA}_{\mathrm{orig}}$) captures semantic foreground structure but provides limited information about actual degradations. The positive and negative prompt attention maps ($\mathrm{CA}_{\mathrm{pos}}$, $\mathrm{CA}_{\mathrm{neg}}$) reveal complementary patterns: $\mathrm{CA}_{\mathrm{pos}}$ tends to emphasize well-formed regions aligned with the prompt, while $\mathrm{CA}_{\mathrm{neg}}$ highlights visually inconsistent or low-confidence areas. Taking their contrast produces $\mathrm{CA}_{\mathrm{diff}}^{\prime}$, which reliably exposes localized artifacts such as missing limbs, distorted geometry, texture noise, or incorrect shading.
Next, the propagated attention $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ spreads semantic information through self-attention layers, sharpening object boundaries and reinforcing regions that are coherent with the global context. When combined ($\mathrm{CA}_{\mathrm{diff}}^{\prime}+\mathrm{CA}_{\mathrm{orig}}^{\prime}$), the resulting mask becomes both defect-aware and semantically grounded: irrelevant background areas are suppressed, while problematic regions remain dominant.
As shown in Fig. 1, SD2.1 masks tend to be smoother and highlight larger defective zones. In contrast, FLUX (Fig. 2) produces finer-grained and more spatially precise masks, reflecting its higher visual capacity and ability to reveal subtle local degradations (e.g., thin structure distortions, glass reflections, or geometric irregularities). Across diverse prompts (human bodies, animals, food, vehicles, and indoor scenes) LoTTS consistently localizes error-prone regions, demonstrating its robustness.
Figure 2: Additional examples of mask generation in FLUX. LoTTS produces finer and more detailed masks in FLUX, consistently localizing local degradations under varied prompts and enabling more precise localized resampling.
@@ C.2 Localized Refinement Examples
Appendix D Ablations
D.1 Mask Generation
Random mask: too fragmented and semantically inconsistent.   The second-worst performance comes from using a random mask. Unlike the “w/o Mask” baseline, the random mask still enforces locality, but it does so in an unstructured and highly fragmented manner. Because foreground regions and background pixels are randomly mixed, the mask fails to form coherent spatial blocks, making refinement unstable and preventing the model from effectively correcting artifacts. This produces masks with both low precision and low recall, explaining why the random baseline consistently underperforms.
Removing Propagation: masks lack spatial continuity.   The best-performing ablation variant is “w/o Propagation.” The mask still identifies the correct semantic and defective regions thanks to both cross-attention terms, but lacks spatial smoothness. Consequently, the mask is less contiguous and refines regions in a slightly noisy manner. This produces moderate performance drops, indicating that while Propagation is not as critical as the attention signals, it contributes substantially to stable, coherent refinement.
Summary.   Across all variants, the degradation trend follows:
Table 3: Ablation study of mask generation strategies on Pick-a-Pic and DrawBench. Results for SD2.1, SDXL, and FLUX show that removing or altering components (e.g., w/o Mask, Random Mask, w/o $\mathrm{CA}_{\mathrm{orig}}^{\prime}$, w/o $\mathrm{CA}_{\mathrm{diff}}$, or w/o Propagation) degrades performance, confirming the effectiveness of our full design.
@@ D.2 Prompt Design
As in Tab. 4, this ablation investigates how different prompt formulations influence the effectiveness of LoTTS. We compare two strategies: (1) “a good/bad image of {p}”, and (2) “a high/low-quality image of {p}”, where {p} denotes the original text prompt. These two designs differ subtly in semantics: the former focuses on the overall desirability of the image, while the latter emphasizes rendering fidelity and fine-grained visual quality. Across all three model families (SD2.1, SDXL, and FLUX), both prompt strategies consistently improve generation quality, with only mild differences between them. For SD2.1 and FLUX, the two prompts perform similarly across all metrics, with the “high/low-quality” formulation yielding slightly higher Pick and IR scores. For SDXL, however, the “good/bad” prompt generally achieves better HPS and AES on Pick-a-Pic, suggesting that SDXL is more responsive to value-oriented descriptors than to explicit quality modifiers. Overall, the results indicate that LoTTS is robust to the specific choice of prompt template: both designs effectively provide the contrastive supervision needed for preference-aware optimization. The observed differences are small and model-specific, implying mild lexical sensitivity but no strong dependence on prompt wording.
@@ E.1 Hyperparameter Sensitivity
Mask area ratio $r$.   Tab. 6 shows that a moderate mask area ratio (about $50\%$) consistently yields the best performance across SD2.1, SDXL, and FLUX. This trend reflects a fundamental trade-off between defect coverage and structural preservation. When the ratio is small (e.g., $r=20\%$), the mask typically captures only the most salient defect regions while ignoring secondary or subtle artifacts. As a result, many imperfections remain unrefined, lowering both the improvement rate (IR) and aesthetic gains (HPS, AES). This under-refinement behavior is evident in SD2.1, SDXL, and FLUX, where $r=20\%$ systematically underperforms the $50\%$ setting across nearly all metrics. Increasing the ratio to a moderate level ($r=50\%$) substantially improves all metrics. At this setting, the mask covers not only primary defects but also neighboring regions that contribute to local inconsistency, enabling more comprehensive correction. This leads to strong gains in HPS, AES, and IR across all three models. However, when the mask becomes too large ($r=80\%$), performance begins to decline. Excessively large masks start to overwrite high- quality regions, particularly detailed textures or well-formed structures, making the refinement process resemble global resampling. Interestingly, FLUX obtains the highest AES at $r=80\%$, but its HPS, Pick, and IR metrics do not surpass the $50\%$ setting, indicating that although aesthetic smoothness may improve, semantic fidelity and user preference do not benefit from excessive mask coverage. Overall, the optimal range around $50\%$ aligns with the observation that localized refinement should balance two competing goals: (1) covering enough area to correct both major and subtle defects, while (2) preserving globally coherent structure and avoiding unnecessary alteration of high-quality content. This trade-off consistently favors a moderate ratio across all tested models and datasets.
@@
Noise injection step $t_{0}$.   Tab. 7 shows that the optimal noise injection step varies noticeably across SD2.1, SDXL, and FLUX, reflecting the different generative dynamics of their underlying architectures. For the U-Net–based diffusion models SD2.1 and SDXL, injecting noise at an early-to-mid denoising step produces the strongest results. Specifically, SD2.1 achieves its best performance at $t_{0}=25$ out of 50 steps, while SDXL peaks at $t_{0}=15$ out of 30 steps. At these intermediate stages, the model retains enough semantic structure from the original sample to preserve the global layout while still having sufficient freedom to update defect regions. This balance leads to clear gains in HPS, AES, and IR. The curves in Tab. 7 further show that both too-early injection ($t_{0}=0$) and too-late injection ($t_{0}=40$ for SD2.1, $t_{0}=25$ for SDXL) result in inferior performance: injecting noise too early disrupts global structure, while injecting too late limits the model’s ability to make meaningful local corrections. In contrast, FLUX exhibits a markedly different pattern due to its rectified flow architecture. Unlike U-Net diffusion, FLUX rapidly forms stable global structures at the beginning of its generation trajectory. As a result, even an extremely small noise step such as $t_{0}=1$ provides enough flexibility to refine local defects without disturbing the overall composition. This is reflected in its consistently strong performance at $t_{0}=1$, achieving the best HPS, AES, Pick, and IR across both datasets. Larger injection steps, however, degrade performance: injecting noise at $t_{0}=5$ or $8$ overwrites too much of the already-converged global structure, causing the refined image to deviate from the intended semantics. Taken together, these observations highlight that the optimal noise injection stage is tightly coupled to the generative process of each architecture. U-Net–based diffusion models favor early-to-mid stages where both structure and flexibility coexist, whereas rectified flow models like FLUX require only a minimal noise injection due to their fast convergence properties. LoTTS naturally adapts to these dynamics by allowing $t_{0}$ to control the balance between preserving global coherence and enabling local refinement.

### v3

@@ B.3 Localized Resampling Procedure
1:Input image $x_{\mathrm{in}}$ (at t=0t{=}0), prompt $p$, total steps $T$, re-noise step $t_{0}$, integration step $t_{g}$, step size $\Delta t$, mask $M$, diffusion model, verifier $V$
@@
4:Global re-noise to $t_{0}$
5:Sample $z_{\text{bg}},z_{\text{mask}}\sim\mathcal{N}(0,I)$
@@
7:Masked refinement ($t_{0}\rightarrow t_{g}$)
8:for $t=t_{0},\,t_{0}-\Delta t,\,\dots,\,t_{g}$ do
9:  Sample $z_{t}\sim\mathcal{N}(0,I)$
10:  $x_{t-\Delta t}\leftarrow(1-M)\odot(\alpha(t{-}\Delta t)x_{\text{anchor}}+\sigma(t{-}\Delta t)z_{t})+M\odot(x_{t}-\Delta t\,\epsilon_{\theta}(x_{t},t;p)+\sigma(t{-}\Delta t)z_{t})$
11:end for
12:Global integration ($t_{g}\rightarrow 0$)
13:for $t=t_{g},\,t_{g}-\Delta t,\,\dots,\,0$ do
14:  Sample $z_{t}\sim\mathcal{N}(0,I)$
15:  $x_{t-\Delta t}\leftarrow x_{t}-\Delta t\,\epsilon_{\theta}(x_{t},t;p)+\sigma(t)z_{t}$
16:end for
@@
19:return $x_{\mathrm{out}},\,v_{\mathrm{out}}$
20:Note: $T$ is included for scheduler compatibility; discretization uses $\Delta t$.
@@ B.4 Overall Algorithm of LoTTS
1:Prompt $p$, DFS depth D=2D{=}2, global seeds $S$, localized refinements $k$, total steps $T$, re-noise $t_{0}$, integration $t_{g}$, step size $\Delta t$, verifier $V$
@@
22:   for $j=1$ to $k$ do
23:     $(x_{\mathrm{out}},v_{\mathrm{out}})\leftarrow\text{LocalizedResample}(x,p,T,t_{0},t_{g},\Delta t,M,V)$
@@ C.1 Mask Generation Examples
Fig. 12, Fig. 13, and Fig. 14 provide a detailed visualization of the mask generation process in SD2.1, SDXL, and FLUX. Across all models, LoTTS exhibits several consistent behaviors that highlight its ability to identify and isolate low-quality regions.
Figure 13: Additional examples of mask generation in SDXL. Compared to SD2.1, LoTTS produces sharper and more spatially precise masks in SDXL, accurately delineating defect boundaries for targeted refinement.
First, the original cross-attention map ($\mathrm{CA}_{\mathrm{orig}}$) captures semantic foreground structure but provides limited information about actual degradations. The positive and negative prompt attention maps ($\mathrm{CA}_{\mathrm{pos}}$, $\mathrm{CA}_{\mathrm{neg}}$) reveal complementary patterns: $\mathrm{CA}_{\mathrm{pos}}$ tends to emphasize well-formed regions aligned with the prompt, while $\mathrm{CA}_{\mathrm{neg}}$ highlights visually inconsistent or low-confidence areas. Taking their contrast produces $\mathrm{CA}_{\mathrm{diff}}$, which reliably exposes localized artifacts such as missing limbs, distorted geometry, texture noise, or incorrect shading.
Next, the propagated attention $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ spreads semantic information through self-attention layers, sharpening object boundaries and reinforcing regions that are coherent with the global context. When combined ($\mathrm{CA}_{\mathrm{diff}}^{\prime}+\lambda\,\mathrm{CA}_{\mathrm{orig}}^{\prime}$), the resulting mask becomes both defect-aware and semantically grounded: irrelevant background areas are suppressed, while problematic regions remain dominant.
As shown in Fig. 12, SD2.1 masks tend to be smoother and highlight larger defective zones. SDXL (Fig. 13) operates at $1024{\times}1024$ resolution, producing masks that are noticeably sharper and more spatially precise than SD2.1, accurately delineating defect boundaries while retaining strong semantic grounding. FLUX (Fig. 14) pushes this further, producing the finest-grained masks that capture subtle local degradations (e.g., thin structure distortions, glass reflections, or geometric irregularities), reflecting its higher visual capacity. Across diverse prompts (human bodies, animals, food, vehicles, and indoor scenes), LoTTS consistently localizes error-prone regions, demonstrating its robustness.
Figure 14: Additional examples of mask generation in FLUX. LoTTS produces finer and more detailed masks in FLUX, consistently localizing subtle degradations under varied prompts and enabling more precise localized resampling.
@@ C.2 Localized Refinement Examples
Table 5: Ablation study of mask generation strategies on Pick-a-Pic and DrawBench. Results for SD2.1, SDXL, and FLUX show that removing or altering components (e.g., w/o Mask, Random Mask, w/o $\mathrm{CA}_{\mathrm{orig}}^{\prime}$, w/o $\mathrm{CA}_{\mathrm{diff}}^{\prime}$, or w/o Propagation) degrades performance, confirming the effectiveness of our full design.
@@
Appendix D Ablations
D.1 Mask Generation
Random mask: too fragmented and semantically inconsistent. The second-worst performance comes from using a random mask. Unlike the “w/o Mask” baseline, the random mask still enforces locality, but it does so in an unstructured and highly fragmented manner. Because foreground and background regions are randomly mixed, the mask fails to form coherent spatial blocks, making refinement unstable and preventing the model from effectively correcting artifacts. This produces masks with both low precision and low recall, explaining why the random baseline consistently underperforms.
Removing Propagation: masks lose spatial continuity. The best-performing ablation variant is “w/o Propagation”. The mask still captures the correct semantic and defective regions due to the two cross-attention signals, but lacks spatial smoothness. As a result, the mask becomes fragmented and refinement occurs in slightly noisy regions. Performance drops moderately, indicating that Propagation improves the spatial stability of the refinement.
Summary. Across all variants, the degradation trend follows:
@@ D.2 Prompt Design
As shown in Tab. 6, this ablation investigates how different prompt formulations influence the effectiveness of LoTTS. We compare two strategies: (1) “A good/bad image of {p}”, and (2) “A high/low-quality image of {p}”, where {p} denotes the original text prompt. These two designs differ subtly in semantics: the former focuses on the overall desirability of the image, while the latter emphasizes rendering fidelity and fine-grained visual quality. Across all three model families (SD2.1, SDXL, and FLUX), both prompt strategies consistently improve generation quality, with only mild differences between them. For SD2.1 and FLUX, the two prompts perform similarly across all metrics, with the “high/low-quality” formulation yielding slightly higher Pick and IR scores. For SDXL, however, the “good/bad” prompt generally achieves better HPS and AES on Pick-a-Pic, suggesting that SDXL is more responsive to value-oriented descriptors than to explicit quality modifiers. Overall, the results indicate that LoTTS is robust to the specific choice of prompt template: both designs effectively provide the contrastive supervision needed for preference-aware optimization. The observed differences are small and model-specific, implying mild lexical sensitivity but no strong dependence on prompt wording.
@@
D.3 Attention Reweighting Coefficient $\lambda$
We ablate the semantic-guided reweighting coefficient $\lambda$ in Eq. equation 7, which balances the contrastive defect signal $\mathrm{CA}^{\prime}_{\mathrm{diff}}$ and the original foreground prior $\mathrm{CA}^{\prime}_{\mathrm{orig}}$. Tab. 7 reports results across SD2.1, SDXL, and FLUX on Pick-a-Pic and DrawBench.
When $\lambda{=}0$, the mask relies solely on the contrastive signal and loses its semantic foreground constraint, often activating on background regions. This is functionally equivalent to the “w/o $\mathrm{CA}_{\mathrm{orig}}^{\prime}$” variant in Tab. 5 and yields the weakest performance across all models and metrics. Performance peaks at $\lambda{=}0.5$ across all three models, striking the best balance between defect sensitivity and foreground grounding. When $\lambda$ is further increased to $1.0$, the foreground prior begins to dominate, causing the mask to spread across the entire foreground rather than focusing on defective regions. This leads to a modest but consistent decline: for example, SD2.1 HPS on Pick-a-Pic drops from 24.52 ($\lambda{=}0.5$) to 24.01 ($\lambda{=}1.0$), while SDXL IR on DrawBench decreases from 1.111 to 1.055. Notably, $\lambda{=}1.0$ still significantly outperforms $\lambda{=}0$ (e.g., SD2.1 HPS: 24.01 vs. 20.43), confirming that even an over-weighted foreground prior is far more beneficial than having none at all. Overall, the results demonstrate that LoTTS is robust to the choice of $\lambda$, with the degradation from the optimal $\lambda{=}0.5$ to $\lambda{=}1.0$ being much smaller than the gap between $\lambda{=}0$ and $\lambda{=}0.5$.
Table 7: Ablation on attention reweighting coefficient $\lambda$. Results on Pick-a-Pic and DrawBench for SD2.1, SDXL, and FLUX.
Model $\lambda$ Pick-a-Pic DrawBench
[…]
SD2.1 0 20.43 5.223 20.87 0.444 21.22 5.570 20.44 0.445
0.50 24.52 5.805 21.32 0.680 23.29 5.911 21.47 0.698
1.00 24.01 5.799 21.29 0.643 23.10 5.835 21.33 0.655
SDXL 0 24.00 6.100 22.00 0.748 24.55 6.222 22.28 0.766
0.50 28.23 6.304 22.30 1.102 28.90 6.321 22.38 1.111
1.00 27.56 6.299 22.22 1.023 26.55 6.300 22.30 1.055
FLUX 0 30.00 6.131 22.45 1.200 30.13 6.220 22.17 1.333
0.50 33.33 6.501 23.04 1.605 33.90 6.890 23.21 1.623
1.00 32.66 6.449 22.89 1.599 33.01 6.655 23.01 1.578
D.4 Consistency Components
We ablate the three consistency mechanisms introduced in Sec. 4 to evaluate their individual contributions to refinement quality. Tab. 8 reports results for SD2.1, SDXL, and FLUX on Pick-a-Pic and DrawBench.
Temporal Consistency (largest degradation). Removing temporal consistency means starting the refinement from pure noise (t=Tt{=}T) rather than the intermediate step $t_{0}$. This forces the model to regenerate the masked region from scratch, destroying the global semantic context and causing the refined content to be semantically inconsistent with the surrounding image. This variant shows the largest degradation across all models: for example, SD2.1 HPS drops from 24.52 to 20.33 on Pick-a-Pic and IR drops from 0.680 to 0.235, falling back to the level of a single unrefined generation. The pattern is consistent for SDXL and FLUX, confirming that preserving global structure via partial denoising is critical.
Spatial Consistency (substantial degradation). Removing spatial consistency means injecting noise only within the mask while keeping the background latent clean at step $t_{0}$. This creates a discontinuous noise distribution at mask boundaries, producing visible seams and edge artifacts. The resulting quality drop is substantial across all metrics, with HPS and IR suffering the most (e.g., SDXL HPS drops from 28.23 to 24.99, IR drops from 1.102 to 0.894 on Pick-a-Pic), as boundary artifacts degrade both perceptual preference and overall image quality.
Holistic Consistency (moderate degradation). Removing holistic consistency means skipping the global integration phase (setting $t_{g}{=}0$, i.e., no unmasked denoising steps after refinement). Without these final harmonization steps, subtle boundary inconsistencies in lighting, color, and texture persist. The degradation is the smallest among the three components but remains consistent: for instance, SD2.1 IR drops from 0.680 to 0.650 on Pick-a-Pic. This demonstrates that even a few global denoising steps effectively smooth residual artifacts and contribute to final quality.
Table 8: Ablation on consistency components. Results on Pick-a-Pic and DrawBench for SD2.1, SDXL, and FLUX.
Model Variant Pick-a-Pic DrawBench
[…]
SD2.1 w/o Temporal 20.33 5.298 20.24 0.235 21.32 5.423 20.21 0.240
w/o Spatial 21.36 5.390 20.69 0.388 21.93 5.584 20.58 0.288
w/o Holistic 24.24 5.799 21.22 0.650 23.09 5.899 21.40 0.687
Ours 24.52 5.805 21.32 0.680 23.29 5.911 21.47 0.698
SDXL w/o Temporal 24.21 6.178 21.89 0.789 25.01 6.220 22.01 0.742
w/o Spatial 24.99 6.225 21.94 0.894 25.55 6.256 22.45 0.789
w/o Holistic 28.05 6.299 22.18 1.008 28.59 6.305 22.35 1.105
Ours 28.23 6.304 22.30 1.102 28.90 6.321 22.38 1.111
FLUX w/o Temporal 30.12 6.288 22.01 1.030 29.19 6.215 22.02 1.084
w/o Spatial 30.89 6.301 22.33 1.088 29.58 6.225 22.56 1.146
w/o Holistic 33.01 6.488 23.03 1.599 33.49 6.823 23.11 1.598
Ours 33.33 6.501 23.04 1.605 33.90 6.890 23.21 1.623
@@ E.1 Hyperparameter Sensitivity
Mask area ratio $r$. Tab. 10 shows that a moderate mask area ratio (about $50\%$) consistently yields the best performance across SD2.1, SDXL, and FLUX. This trend reflects a fundamental trade-off between defect coverage and structural preservation. When the ratio is small (e.g., $r=20\%$), the mask typically captures only the most salient defect regions while ignoring secondary or subtle artifacts. As a result, many imperfections remain unrefined, lowering both the improvement rate (IR) and aesthetic gains (HPS, AES). This under-refinement behavior is evident in SD2.1, SDXL, and FLUX, where $r=20\%$ systematically underperforms the $50\%$ setting across nearly all metrics. Increasing the ratio to a moderate level ($r=50\%$) substantially improves all metrics. At this setting, the mask covers not only primary defects but also neighboring regions that contribute to local inconsistency, enabling more comprehensive correction. This leads to strong gains in HPS, AES, and IR across all three models. However, when the mask becomes too large ($r=80\%$), performance begins to decline. Excessively large masks start to overwrite high- quality regions, particularly detailed textures or well-formed structures, making the refinement process resemble global resampling. Interestingly, on Pick-a-Pic, FLUX obtains the highest AES at $r=80\%$, but its HPS, Pick, and IR metrics do not surpass the $50\%$ setting, indicating that although aesthetic smoothness may improve, semantic fidelity and user preference do not benefit from excessive mask coverage. Overall, the optimal range around $50\%$ aligns with the observation that localized refinement should balance two competing goals: (1) covering enough area to correct both major and subtle defects, while (2) preserving globally coherent structure and avoiding unnecessary alteration of high-quality content. This trade-off consistently favors a moderate ratio across all tested models and datasets.
@@
Noise injection step $t_{0}$. Tab. 11 shows that the optimal noise injection step varies noticeably across SD2.1, SDXL, and FLUX, reflecting the different generative dynamics of their underlying architectures. For the U-Net–based diffusion models SD2.1 and SDXL, injecting noise at an early-to-mid denoising step produces the strongest results. Specifically, SD2.1 achieves its best performance at $t_{0}=25$ out of 50 steps, while SDXL peaks at $t_{0}=15$ out of 30 steps. At these intermediate stages, the model retains enough semantic structure from the original sample to preserve the global layout while still having sufficient freedom to update defect regions. This balance leads to clear gains in HPS, AES, and IR. The results in Tab. 11 further show that both too-early injection ($t_{0}=0$) and too-late injection ($t_{0}=40$ for SD2.1, $t_{0}=25$ for SDXL) result in inferior performance: injecting noise too early limits the model’s ability to make meaningful local corrections, while injecting too late disrupts global structure. In contrast, FLUX exhibits a markedly different pattern due to its rectified flow architecture. Unlike U-Net diffusion, FLUX rapidly forms stable global structures at the beginning of its generation trajectory. As a result, even an extremely small noise step such as $t_{0}=1$ provides enough flexibility to refine local defects without disturbing the overall composition. This is reflected in its consistently strong performance at $t_{0}=1$, achieving the best HPS, AES, Pick, and IR across both datasets. Larger injection steps, however, degrade performance: injecting noise at $t_{0}=5$ or $8$ overwrites too much of the already-converged global structure, causing the refined image to deviate from the intended semantics. Taken together, these observations highlight that the optimal noise injection stage is tightly coupled to the generative process of each architecture. U-Net–based diffusion models favor early-to-mid stages where both structure and flexibility coexist, whereas rectified flow models like FLUX require only a minimal noise injection due to their fast convergence properties. LoTTS naturally adapts to these dynamics by allowing $t_{0}$ to control the balance between preserving global coherence and enabling local refinement.